# Beyond Chunking: Discourse-Aware Hierarchical Retrieval for Long Document Question Answering

### v3

@@ Abstract
Long document question answering systems typically process texts as flat sequences or use arbitrary segmentation, failing to capture discourse structures that guide human comprehension. We present a discourse-aware hierarchical framework that leverages rhetorical structure theory (RST) to enhance long document question answering. Our approach converts discourse trees into sentence-level representations and employs LLM-enhanced node representations to bridge structural and semantic information. The framework involves three key innovations: specialized discourse parsing for lengthy documents, LLM-based enhancement of discourse relation nodes, and structure-guided hierarchical retrieval. Comprehensive experiments on QASPER, QuALITY, and NarrativeQA demonstrate consistent improvements over existing approaches. Ablation studies confirm that incorporating discourse structure significantly enhances question answering across diverse document types.
@@
Huiyao Chen,   Yi Yang,   Yinghui Li,  Meishan Zhangthanks: Corresponding Author: Meishan Zhang,   Min Zhang Institute of Computing and Intelligence, Harbin Institute of Technology (Shenzhen), China {chenhy1018,yangi10010730,mason.zms}@gmail.com
@@ 1 Introduction
Therefore, in this work, we present DISRetrieval (DIScourse-aware hierarchical Retrieval), the first systematic approach that leverages rhetorical structure theory (RST) to enhance long document question answering through discourse-aware modeling. Our approach tackles three key technical challenges: First, we develop an effective adaptation strategy that enables traditional RST to work at appropriate granularity levels for computational efficiency while preserving semantic coherence. Second, we introduce LLM-based node enhancement that enriches intermediate nodes with both discourse structure and semantic content. Third, we design structure-guided evidence retrieval mechanisms that leverage discourse organization for more effective information extraction.
Comprehensive experiments across three challenging benchmarks: QASPER (Dasigi et al., 2021) for research paper understanding, QuALITY (Pang et al., 2022) for reading comprehension, and NarrativeQA Kočiský et al. (2018) for book-length narrative analysis demonstrate substantial improvements over existing methods. Detailed ablation studies validate the effectiveness of our discourse-aware modeling and structure-guided processing, showing consistent gains across diverse scenarios. Our framework successfully captures both fine-grained semantic details and document-level organizational patterns, providing a robust solution for discourse-informed question answering.
Our main contributions can be summarized as follows:
  • A novel discourse-aware hierarchical framework that systematically incorporates RST for long document question answering.
  • An innovative LLM-enhanced hierarchical retrieval mechanism that enriches intermediate nodes with semantic content and enables multi-granularity retrieval.
  • Comprehensive empirical validation demonstrating consistent improvements across diverse datasets and model architectures.
@@
2 Related Works
Recent advances in natural language processing have achieved remarkable progress in document processing and understanding, with large language models demonstrating impressive capabilities across diverse text comprehension tasks (Brown et al., 2020; Chowdhery et al., 2023). However, these successes primarily focus on shorter texts, as extended documents present fundamental challenges due to context length constraints and computational complexity (Tay et al., 2021; Zaheer et al., 2020; Beltagy et al., 2020; Ding et al., 2023; Ainslie et al., 2023; Chen et al., 2024). Traditional approaches typically operate on short text segments without considering broader document context, leading to fragmented understanding (Liu and Lapata, 2019; Guo et al., 2022; Rae et al., 2020). Recent innovations such as SEER’s chunking-free key information extraction (Zhao et al., 2024) and CFIC’s in-context retrieval approach (Qian et al., 2024) highlight the ongoing evolution in document comprehension strategies for extended content (Wang et al., 2023; Izacard et al., 2023; Yu et al., 2023).
To tackle these challenges, retrieval-based methods have emerged as a promising paradigm, segmenting documents into manageable pieces before retrieving the most relevant evidence for downstream tasks (Lewis et al., 2020; Karpukhin et al., 2020; Guu et al., 2020a; Izacard and Grave, 2020). The evolution of retrieval-based approaches has progressed from simple flat segmentation strategies (Guu et al., 2020b) to sophisticated hierarchical organization methods. RAPTOR pioneered bottom-up semantic clustering for tree-based document representation (Sarthi et al., 2024), while bisection-based methods preserve local coherence through adjacent segment grouping (Zhang et al., 2020; Ivgi et al., 2023). Recent innovations include hybrid hierarchical retrieval systems that bridge document-level and passage-level representations (Liu et al., 2021; Arivazhagan et al., 2023), and LongRefiner’s XML-format structure conversion using fine-tuned LLMs for top-down tree construction, though its length sensitivity creates challenges for long documents (Jin et al., 2025). While these hierarchical approaches represent significant progress in document modeling, they primarily rely on surface-level semantic similarity or structural heuristics, overlooking the rich linguistically-informed discourse structures that naturally organize human-authored documents (Liu et al., 2019; Wolf et al., 2020).
@@ 3 Method
Discourse analysis provides systematic frameworks for understanding how textual segments relate beyond surface-level semantics, offering principled approaches to document organization Mann and Thompson (1988); Carlson et al. (2001). RST formalizes this understanding by representing documents as hierarchical trees where leaf nodes correspond to elementary discourse units (EDUs) and internal nodes encode rhetorical relations such as contrast, elaboration, and summary Marcu (2000). Recent advances in neural discourse parsing Yu et al. (2018); Kobayashi et al. (2020); Yu et al. (2022a) have enabled automatic RST construction at scale, facilitating its application across diverse NLP tasks Maekawa et al. (2024); Yuan et al. (2025).
To address the challenge of bridging between discourse theory and document retrieval, we propose DISRetrieval, a discourse-aware framework that systematically incorporates RST structure into document retrieval as shown in Figure 2. The complete algorithm is detailed in Algorithm 2 in Appendix A.
Figure 2: Overview of the DISRetrieval framework. The framework consists of three main steps: (1) Discourse-Aware Tree Construction that builds a hierarchical discourse structure through two phases: constructing paragraph-level discourse trees via discourse parsing, and generating document-level tree with LLM-enhanced node representations; (2) Node Representation that integrates the trees and converts text content into dense vector representations via an encoder; (3) Hierarchical Evidence Retrieval and Selection that performs multi-level evidence retrieval and structure-aware selection to identify relevant text segments.
@@ 3.1 Discourse-Aware Tree Construction
The first stage of our framework aims to construct a hierarchical tree structure that captures semantic content and organizational structure of documents. We leverage discourse analysis to identify meaningful relationships between text segments and organize them into a coherent tree representation.
3.1.1 Granularity-Adapted RST Processing
Traditional RST operates on fine-grained EDUs, which can create computational overhead and semantic fragmentation when processing long documents. We develop a granularity adaptation strategy that shifts RST processing to the sentence level, achieving better semantic coherence and efficiency while maintaining discourse interpretability. Our adaptation involves training a sentence-level discourse parser by systematically converting EDU-based datasets: we merge intra-sentence EDUs into coherent sentence units and determine inter-sentence relationships through lowest common ancestor analysis in the original EDU trees. This produces a specialized parser $f_{\text{discourse}}$ that operates directly on sentence-level discourse relationships, enabling efficient processing of long documents.
@@ 3.1.2 Complete Tree Construction Process
Our core innovation lies in developing a hierarchical discourse modeling framework that constructs complete document-level discourse trees through a systematic two-phase process, effectively bridging local sentence-level relationships with global document organization while preserving discourse coherence at multiple granularity levels.
@@
The initial phase processes each paragraph $p_{i}$ by transforming its sentence sequence $\bm{S_{i}}=\{s_{i,1},s_{i,2},...,s_{i,m}\}$ into a local discourse tree:
$T_{i}=f_{\text{discourse}}(\bm{S_{i}}),$ (1)
where each tree $T_{i}$ captures how sentences connect and build upon each other within the paragraph, reflecting both semantic relationships and logical flow between adjacent sentences. This phase ensures that fine-grained intra-paragraph discourse structure is preserved.
@@
The second phase constructs the global document structure by first obtaining simplified representations of paragraphs through bottom-up LLM enhancement. For each internal node $v$ in paragraph-level trees, we apply adaptive processing:
where $v_{l}$ and $v_{r}$ denote the left and right children of node $v$. The function $f_{\text{LLM}}$ generates concise summaries for larger text spans that align with our sentence-level discourse parser, while $f_{\text{merge}}(v)$ merges smaller, coherent subtrees into single leaf nodes to maintain semantic coherence.
The root representations obtained from each paragraph tree serve as input for constructing the document-level discourse tree:
$T_{\text{doc}}=f_{\text{discourse}}(\{t_{\text{root}}^{1},t_{\text{root}}^{2},...,t_{\text{root}}^{n}\}).$ (3)
This approach transforms rhetorical relations into concrete semantic summaries that effectively represent the content at different levels of the discourse hierarchy (detailed illustration provided in Appendix A.3).
@@ 3.2 Discourse-Aware Node Representation
A critical challenge in discourse-based retrieval lies in bridging the semantic gap between abstract rhetorical relations and concrete textual content required for neural retrieval. We develop a semantic enhancement framework that transforms discourse structures into dense representations while preserving hierarchical properties through three key stages.
@@
The document-level discourse tree $T_{\text{doc}}$ initially contains only abstract structural relationships at internal nodes, lacking concrete semantic content necessary for retrieval. We apply bottom-up LLM enhancement following Equation 2 to generate meaningful textual representations for these structural nodes. This enhancement process systematically transforms each internal node from abstract structural placeholders into semantically rich representations that capture the combined content of their subtrees, ensuring all nodes possess concrete textual content suitable for downstream semantic matching operations.
@@ 3.3 Structure-Guided Evidence Retrieval
Direct leaf selection: High-relevance sentences selected immediately Hierarchical expansion: Internal nodes trigger top-$k$ subtree selection Redundancy elimination: Automatic tracking prevents duplicate selection
@@
The complete algorithm is detailed in Algorithm 1, which naturally balances local semantic relevance with global discourse coherence, enabling retrieval at appropriate granularity levels determined by the document’s discourse structure.
@@ 4.1 Experimental Setup
We evaluate our approach on three challenging long document QA datasets: QASPER (Dasigi et al., 2021) for research paper understanding (averaging 4170 words), QuALITY (Pang et al., 2022) for long-form reading comprehension (averaging 5022 words), and NarrativeQA (Kočiský et al., 2018) for QA over exceptionally long narrative documents (averaging 51,372 words). For QASPER, we use F1-Match for QA performance and token-level F1/Recall for retrieval evaluation. For QuALITY, we report accuracy as the primary metric. And for NarrativeQA, we measure performance using BLEU(B-1), ROUGE(R-L) and METEOR(M) metrics. To ensure fair comparison, we fix the total length of retrieved context across all methods.
@@
  • Bisection shares our LLM-enhanced representations and hierarchical retrieval mechanism but uses a different tree construction method - recursively dividing sentence set into balanced binary subtrees. This baseline helps isolate benefits of our discourse-aware tree structure.
@@
For discourse analysis, we construct our sentence-level RST training data based on RST-DT (Carlson et al., 2001). We then train our discourse parser following Yu et al. (2022b), adopting XLNet-base-cased as the pre-trained language model backbone. All summarization tasks, including both our approach (as defined in Equation 2) and RAPTOR baseline, utilize Llama3.1-8B-Instruct for consistency. We set the threshold parameter $\tau$ in Equation 2 to 0 for QASPER and 50 for QuALITY and NarrativeQA (detailed in Appendix C.4). For sentence encoder (Section §3.2), we experiment with both Sentence-BERT111multi-qa-mpnet-base-cos-v1, https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 and the latest embedding model of OpenAI [^2] OpenAI text-embedding-3-large to evaluate the effectiveness of our method across different representation spaces. The Top-K parameter in leaf selection is set to 5 based on preliminary experiments (detailed in Appendix C.3). For generation models, we employ UnifiedQA-3B, GPT-4.1-mini and Deepseek-v3 to demonstrate the robustness of our retrieval improvements across different generation capabilities. More details are provided in Appendix B.
@@ 4.2 Main Results
    DISRetrieval 38.34 40.03 40.33 42.10 41.61 43.30 68.94 72.96 72.39 76.70 73.78 78.00
Table 1: Generation performance comparison of different methods under varying retrieved context lengths (200-400 words) and different embedding models (SBERT: s, OpenAI: o) across two datasets. Best results are bolded, runner-ups are underlined.
Consistent superiority across settings. DISRetrieval outperforms all baselines regardless of context length, embedding model, or generation architecture (UnifiedQA-3B, GPT-4.1-mini, and Deepseek-v3). For instance, with 400-word contexts and UnifiedQA-3B, we achieve +2.58% F1-Match on QASPER and +4.37% accuracy on QuALITY over the flatten-sentence baseline.
Discourse structure surpasses semantic clustering. Compared to RAPTOR’s semantic clustering approach, our discourse-aware method demonstrates clear advantages (40.24% vs. 39.96% F1-Match on QASPER with OpenAI embeddings), confirming that linguistic discourse structure provides more principled document organization than purely semantic-based methods.
@@
200 (SBERT) 200 (OpenAI) 300 (SBERT) 300 (OpenAI) 400 (SBERT) 400 (OpenAI) Avg. 4170 wrods Avg. 129 words
F1 / % 49.56 (+0.75) 49.83 (+1.02) 49.63 (+0.82) 49.72 (+0.91) 49.71 (+0.90) 49.56 (+0.75) 48.81 50.71 (+1.90)
Table 2: Retrieval and generation results (token-level F1 and Recall) on the QASPER dataset using SBERT and OpenAI Embedding as embedding models and Llama3.1-8B-Instruct as generation model. Best results are bolded and runner-ups are underlined.
@@
DISRetrieval 25.59 30.26 21.21
Table 3: Performance on the NarrativeQA dataset. Best results are bolded and runner-ups are underlined.

### v4

@@ Abstract
Existing long-document question answering systems typically process texts as flat sequences or use heuristic chunking, which overlook the discourse structures that naturally guide human comprehension. We present a discourse-aware hierarchical framework that leverages rhetorical structure theory (RST) for long document question answering. Our approach converts discourse trees into sentence-level representations and employs LLM-enhanced node representations to bridge structural and semantic information. The framework involves three key innovations: language-universal discourse parsing for lengthy documents, LLM-based enhancement of discourse relation nodes, and structure-guided hierarchical retrieval. Extensive experiments on four datasets demonstrate consistent improvements over existing approaches through the incorporation of discourse structure, across multiple genres and languages. Moreover, the proposed framework exhibits strong robustness across diverse document types and linguistic settings.
@@
Huiyao Chen1,2, Yi Yang1, Yinghui Li1, Meishan Zhang1,thanks: Correspondingauthor:MeishanZhang. Baotian Hu1,2, Min Zhang1,2 1Institute of Computing and Intelligence, Harbin Institute of Technology (Shenzhen), China 2 Shenzhen Loop Area Institute (SLAI) {chenhy1018,yangi10010730,mason.zms}@gmail.com
@@ 1 Introduction
Therefore, in this work, we present DISRetrieval (DIScourse-aware hierarchical Retrieval), the first systematic approach that leverages rhetorical structure theory (RST) to enhance long document question answering through discourse-aware modeling with cross-lingual applicability. Our approach tackles three key technical challenges: First, we introduce RST adaptations along two dimensions: granularity adaptation that shifts processing to the sentence level for efficiency, and language adaptation that enables cross-lingual applicability through LLM-based data augmentation. Second, we introduce LLM-based node enhancement that enriches intermediate nodes with both discourse structure and semantic content. Third, we design structure-guided evidence retrieval mechanisms that leverage discourse organization for effective information extraction.
Comprehensive experiments across four challenging benchmarks: QASPER (Dasigi et al., 2021) for research paper understanding, QuALITY (Pang et al., 2022) for reading comprehension, NarrativeQA Kočiský et al. (2018) for book-length narrative analysis and MultiFieldQA-zh (Bai et al., 2024) for Chinese documents question answering demonstrate substantial improvements over existing methods. Detailed ablation studies validate the effectiveness of our discourse-aware modeling and structure-guided processing, showing consistent gains across diverse scenarios. Our framework successfully captures both fine-grained semantic details and document-level organizational patterns, providing a robust solution for discourse-informed question answering. Our main contributions can be summarized as follows:
  • A novel discourse-aware hierarchical framework with granularity and language adaptations enabling efficient cross-lingual long document QA.
  • An innovative LLM-enhanced hierarchical retrieval mechanism enabling multi-granularity evidence selection.
  • Comprehensive empirical validation across diverse datasets, architectures, and languages.
@@
2 Related Work
Recent advances in LLMs have demonstrated impressive capabilities across diverse text comprehension tasks (Brown et al., 2020; Chowdhery et al., 2023). However, extended documents present challenges due to context length constraints and computational complexity (Tay et al., 2021; Zaheer et al., 2020; Beltagy et al., 2020; Ding et al., 2023; Ainslie et al., 2023). Traditional approaches operate on short segments without considering broader context (Liu and Lapata, 2019; Guo et al., 2022; Rae et al., 2020), while recent innovations explore chunking-free extraction (Zhao et al., 2024) and in-context retrieval (Qian et al., 2024; Wang et al., 2023; Izacard et al., 2023; Yu et al., 2023).
Retrieval-based methods segment documents before retrieving relevant evidence (Lewis et al., 2020; Karpukhin et al., 2020; Izacard and Grave, 2020). These have evolved from flat segmentation (Guu et al., 2020b) and early hierarchical approaches (Tang et al., 2017) to sophisticated methods including semantic clustering (Sarthi et al., 2024), bisection-based techniques (Zhang et al., 2020; Ivgi et al., 2023), hybrid systems (Liu et al., 2021; Arivazhagan et al., 2023), and structure conversion approaches (Jin et al., 2025). Discourse structure has been explored for QA through discourse-based systems (Santhosh and Ali, 2012), long-form answer analysis (Xu et al., 2022), and structure-discourse graphs (Nair et al., 2023; Du et al., 2023). However, existing approaches either rely on surface-level semantic similarity for retrieval or apply discourse primarily to answer generation tasks, overlooking linguistically-informed discourse structures for evidence retrieval (Liu et al., 2019; Wolf et al., 2020).
@@ 3 Method
Discourse analysis provides systematic frameworks for understanding how textual segments relate beyond surface-level semantics Mann and Thompson (1988); Carlson et al. (2001). RST formalizes this by representing documents as hierarchical trees where leaf nodes correspond to elementary discourse units (EDUs) and internal nodes encode rhetorical relations such as contrast, elaboration, and summary Marcu (2000). Recent advances in neural discourse parsing Yu et al. (2018); Kobayashi et al. (2020); Yu et al. (2022b, a) have enabled automatic RST construction at scale Maekawa et al. (2024); Yuan et al. (2025).
To address the challenge of bridging between discourse theory and document retrieval, we propose DISRetrieval, a discourse-aware framework that systematically incorporates RST structure into document retrieval as shown in Figure 2. Our framework operates in three stages: (1) constructing hierarchical discourse trees through sentence-level RST parsing, (2) enriching tree nodes with semantic representations via LLM enhancement and dense encoding, and (3) performing structure-guided evidence retrieval. Figure 5 provides a concrete example showing how a research paper paragraph is transformed into a hierarchical discourse structure.
Figure 2: Overview of the DISRetrieval framework consisting of three main steps: (1) Discourse-Aware Tree Construction through paragraph-level parsing (left) and document-level integration (center), (2) Node Representation via tree integration and dense encoding, and (3) Hierarchical Evidence Retrieval with multi-level selection. Color coding indicates discourse relations.
@@ 3.1 Discourse-Aware Tree Construction
We construct a hierarchical tree structure that captures semantic content and organizational structure of documents through discourse analysis.
3.1.1 RST Adaptation
To effectively apply RST to long document question answering, we introduce two critical adaptations that address computational efficiency and cross-lingual applicability.
Granularity Adaptation.
Traditional RST operates on fine-grained EDUs, creating computational overhead and semantic fragmentation for long documents. We address this challenge by shifting RST processing to the sentence level, achieving better efficiency while maintaining semantic coherence. Specifically, we train a sentence-level parser by converting existing EDU-based datasets through two operations: (1) merging intra-sentence EDUs into unified sentence units, and (2) determining inter-sentence relationships via lowest common ancestor analysis in the original EDU-level discourse trees. This enables our parser to capture meaningful discourse relations efficiently.
Language Adaptation.
To enable cross-lingual applicability beyond English, we develop a language-universal discourse parser through LLM-based multilingual data augmentation. We employ GPT-4o to translate the RST-DT training corpus into target languages while preserving discourse structures at the sentence level. The translated data is combined with the original corpus to train a unified parser $f_{\text{discourse}}$ that generalizes across languages without requiring language-specific annotations. Complete implementation details are provided in Appendix A.1 and B.2.
@@ 3.1.2 Complete Tree Construction Process
Building upon our adapted sentence-level discourse parser, we develop a hierarchical discourse modeling framework that constructs document-level discourse trees through a two-phase process.
@@
We transform each paragraph’s sentence sequence $\bm{S_{i}}=\{s_{i,1},s_{i,2},...,s_{i,m}\}$ into a local discourse tree $T_{i}$ using $f_{\text{discourse}}$. Each tree captures sentence connections within the paragraph, preserving fine-grained intra-paragraph discourse structure.
@@
We construct global document structure through bottom-up LLM enhancement. For each internal node $v$ in paragraph-level trees, we apply adaptive processing:
$v^{*}=\begin{cases}f_{\text{LLM}}(v_{l},v_{r}),&\text{if }|v_{l}|+|v_{r}|\geqslant\tau\\ f_{\text{merge}}(v).&\text{otherwise}\end{cases}$ (1)
When combined child length exceeds threshold $\tau$, $f_{\text{LLM}}$ generates concise summaries. Otherwise, $f_{\text{merge}}$ directly concatenates content. Figure 5 illustrates how rhetorical relations transform into semantic representations.
The root representations from paragraph trees serve as input for document-level tree construction. We apply the same RST parser to these summaries, creating a hierarchical structure capturing document-wide discourse relationships.
@@ 3.2 Discourse-Aware Node Representation
A critical challenge in discourse-based retrieval lies in bridging the semantic gap between abstract rhetorical relations and concrete textual content required for neural retrieval. We develop a semantic enhancement framework through three key stages.
@@
The document-level tree $T_{\text{doc}}$ initially lacks concrete semantic content at internal nodes. We apply the same bottom-up LLM enhancement strategy (Equation 1) to generate meaningful textual representations for these structural nodes. Starting from leaf nodes and moving upward, this process systematically transforms each internal node from abstract structural placeholders into semantically rich representations that capture the combined content of their subtrees. The result is a fully enhanced tree where all nodes possess concrete textual content suitable for semantic matching.
@@ 3.3 Structure-Guided Evidence Retrieval
Direct leaf selection for high-relevance sentences Hierarchical expansion with top-$k$ subtree selection for internal nodes Redundancy elimination to prevent duplicates
@@
Algorithm 1 details the complete process, balancing semantic relevance with discourse coherence.
@@ 4.1 Experimental Setup
We evaluate our approach on four challenging long document QA datasets: QASPER (Dasigi et al., 2021) for research papers (avg. 4170 words), QuALITY (Pang et al., 2022) for reading comprehension (avg. 5022 words), NarrativeQA (Kočiský et al., 2018) for narrative documents (avg. 51,372 words), and MultiFieldQA-zh Bai et al. (2024) for Chinese documents (avg. 6701 words). We use F1-Match for QASPER and MultiFieldQA-zh, accuracy for QuALITY, and BLEU(B-1)/ROUGE/METEOR for NarrativeQA. Token-level F1/Recall evaluate retrieval quality on QASPER. We fix the retrieved context length across methods for fair comparison.
@@
  • Bisection shares our LLM-enhanced representations and hierarchical retrieval mechanism (Sections 3.2 and 3.3) but constructs trees by recursively dividing sentences into balanced binary subtrees. This isolates the specific contribution of discourse structure by keeping all other components identical to DISRetrieval.
@@
We train our language-universal discourse parser on RST-DT (Carlson et al., 2001) following Yu et al. (2022a) with gte-multilingual-base111huggingface.co/Alibaba-NLP/gte-multilingual-base backbone Zhang et al. (2024), augmented with GPT-4o-translated Chinese data. All summarization tasks, including both our approach (as defined in Equation 1) and RAPTOR baseline, utilize Llama3.1-8B-Instruct for consistency. We set threshold $\tau=0$ for QASPER and $\tau=50$ for QuALITY/NarrativeQA (Appendix C.4). For sentence encoder (Section §3.2), we test Sentence-BERT222multi-qa-mpnet-base-cos-v1 and OpenAI text-embedding-3-large [^3] https://platform.openai.com/docs/guides/embeddings. Top-K is set to 5 (Appendix C.3). Generation models include UnifiedQA-3B, GPT-4.1-mini, and Deepseek-v3. Details in Appendix B.
@@ 4.2 Main Results
    DISRetrieval 37.79 39.77 40.39 42.19 41.51 43.65 68.89 72.28 72.15 76.46 73.68 77.71
Table 1: Generation performance comparison of different methods under varying retrieved context lengths (200-400 words) and different embedding models (SBERT: s, OpenAI text-embedding-3-large: o) across two datasets. Bisection shares our LLM-enhanced node representations and hierarchical retrieval mechanism but uses binary tree construction instead of discourse structure. Best results are bolded, runners-up are underlined.
Consistent superiority across settings. DISRetrieval outperforms all baselines regardless of context length, embedding model, or generation architecture (UnifiedQA-3B, GPT-4.1-mini, and Deepseek-v3). For instance, with 400-word contexts and UnifiedQA-3B, we achieve +2.66% F1-Match on QASPER and +3.60% accuracy on QuALITY over the flatten-sentence baseline.
Discourse structure surpasses semantic clustering. Compared to RAPTOR’s semantic clustering approach, our discourse-aware method demonstrates clear advantages (40.03% vs. 39.96% F1-Match on QASPER with OpenAI embeddings), confirming that linguistic discourse structure provides more principled document organization than purely semantic-based methods.
@@
Cross-Lingual Effectiveness.
To verify the language-universal capability of our approach, we evaluate on MultiFieldQA-zh, a Chinese long-document QA benchmark. As shown in Table 3, DISRetrieval consistently outperforms all baselines across different context lengths with both GPT-4.1-mini (35.25% F1) and Deepseek-v3 (29.54% F1). The improvements are particularly notable at longer contexts, where our method gains +1.30% to +1.56% over Bisection with Deepseek-v3. These results validate that linguistically-grounded discourse structures transcend language boundaries, with our language-universal parser effectively capturing discourse relationships in Chinese documents. This cross-lingual effectiveness highlights the language-agnostic nature of rhetorical structure theory and demonstrates the extensibility to other languages through similar data augmentation strategies.
@@
200 (SBERT) 200 (OpenAI) 300 (SBERT) 300 (OpenAI) 400 (SBERT) 400 (OpenAI) Avg. 4170 words Avg. 129 words
F1 / % 49.65 (+0.84) 50.05 (+1.24) 49.77 (+0.96) 50.11 (+1.30) 49.54 (+0.73) 50.20 (+1.39) 48.81 50.71 (+1.90)
Table 2: Retrieval and generation results (token-level F1 and Recall) on the QASPER dataset. Embedding models: SBERT and OpenAI text-embedding-3-large. Generation model: Llama3.1-8B-Instruct.
Method GPT-4.1-mini (F1 / %) Deepseek-v3 (F1 / %)
200 300 400 200 300 400
flatten-chunk 28.77 31.25 33.46 22.69 26.91 26.70
flatten-sentence 28.59 31.95 34.41 26.15 27.04 27.06
RAPTOR 26.61 31.70 34.32 23.11 25.96 27.01
Ours
Bisection 28.23 32.31 34.80 26.71 27.05 28.24
DISRetrieval 29.60 32.97 35.25 26.76 28.61 29.54
Table 3: Performance on MultiFieldQA-zh (Chinese).
@@
DISRetrieval 25.39 30.31 21.16
Table 4: Performance on the NarrativeQA dataset.